# Multi-Task-oriented Nighttime Haze Imaging Enhancer for Vision-driven Measurement Systems

### v1

@@ Abstract
Salient object detection (SOD) plays a critical role in vision-driven measurement systems (VMS), facilitating the detection and segmentation of key visual elements in an image. However, adverse imaging conditions such as haze during the day, low light, and haze at night severely degrade image quality, and complicating the SOD process. To address these challenges, we propose a multi-task-oriented nighttime haze imaging enhancer (MToIE), which integrates three tasks: daytime dehazing, low-light enhancement, and nighttime dehazing. The MToIE incorporates two key innovative components: First, the network employs a task-oriented node learning mechanism to handle three specific degradation types: day-time haze, low light, and night-time haze conditions, with an embedded self-attention module enhancing its performance in nighttime imaging. In addition, multi-receptive field enhancement module that efficiently extracts multi-scale features through three parallel depthwise separable convolution branches with different dilation rates, capturing comprehensive spatial information with minimal computational overhead. To ensure optimal image reconstruction quality and visual characteristics, we suggest a hybrid loss function. Extensive experiments on different types of weather/imaging conditions illustrate that MToIE surpasses existing methods, significantly enhancing the accuracy and reliability of vision systems across diverse imaging scenarios. The code is available at https://github.com/Ai-Chen-Lab/MToIE.
Index Terms:
@@
Vision-driven measurement systems, adverse weather, multi-task learning, nighttime haze, image restoration
I Introduction
@@
Salient object detection (SOD) focuses on detecting and segmenting the most visually prominent objects within vision-driven measurement systems (VMS) [1, 2]. It is critical for enhancing automation and accuracy in VMS, particularly in applications like quality control and safety monitoring. However, as shown in Fig. 1, adverse imaging environmental conditions, such as low light, daytime or nighttime haze, can significantly affect the accuracy of SOD. Specifically, optical scattering from haze particles not only reduces contrast and sharpness but also distorts the spectral content of the captured scene [3]. Reduced illumination further exacerbates this issue by limiting signal strength, necessitating longer exposures or higher sensor gains, which result in additional electronic noise [4]. These environmental factors necessitate robust VMS that can effectively model and compensate for light attenuation, atmospheric scattering, and sensor-induced artifacts while maintaining measurement fidelity.
@@
Figure 1: The upper and lower sections illustrate daytime imaging scenarios, where varying haze densities impact object visibility. SOD operates on haze devices as a general-purpose preprocessing module, filtering salient data locally and transmitting it to the cloud for VMS.
@@
To address the challenges posed by harsh imaging conditions, we propose a multi-task-oriented nighttime haze image enhancement framework (MToIE), which be used to handle three types of degradation tasks: ID, LLIE, and NHIE. The framework integrates multi-task learning mechanisms, self-attention (SA) module, and multi-receptive-field enhancement (MRFE) module to effectively enhance image quality under complex degradation scenarios. By employing task-specific node learning mechanisms and effectively using daytime haze, low-light, and nighttime haze datasets for training, MToIE learns common visual patterns (such as edge contours and texture information) across different imaging conditions and specific degradation factors (such as brightness loss and noise enhancement in low-light environments, as well as contrast reduction and scattering blur induced by haze), efficiently models three distinct types of degradation. In addition, a hybrid loss function is designed to balance image reconstruction quality and visual features. The main contributions of this work are summarized as follows
Figure 2: Overall architecture of the multi-task-oriented nighttime haze image enhancement (MToIE) framework. The framework is designed to simultaneously handle three types of degradation tasks: ID, LLIE, and NHIE. By integrating multi-task learning mechanisms, SA module, and MRFE module, it effectively improves image quality under complex degradation scenarios. The whole process of MToIE is summarized in Algorithm 1.
@@
  • We propose a multi-task-oriented nighttime haze imaging enhancer (MToIE) that incorporates task-specific node learning mechanisms and adopts shared and separated networks for feature-shared and specific feature extraction to effectively address low-light, daytime haze, and nighttime haze degradation scenarios.
@@
  • The MToIE integrates SA mechanism and MRFE module to extract multi-scale features. Additionally, a hybrid loss function is suggested to jointly optimize MToIE for robust performance under challenging conditions.
@@
  • Extensive experiments demonstrate that MToIE significantly outperforms existing methods in image quality improvement and achieves higher accuracy and reliability in subsequent high-level vision task.
@@
The remainder of this paper is organized as follows. Degradation models for nighttime haze are given in Section II. MToIE is detailedly described in Section III. Experimental results and discussions are provided in Section IV. Conclusions are drawn in Section V.
II Physical Imaging Models
@@
It needs to be noted that in practical applications of VMS, Eq. 1 needs to be adjusted according to specific circumstances. Due to the complexity of light source, particularly at night, directly applying above formula may not be sufficiently accurate. The challenges posed by varying lighting conditions, haze, and sensor noise require a more customized approach. To address these issues, our method focuses on three key image restoration tasks in VMS: ID, LLIE, and NHIE. To improve model accuracy under these conditions, we propose enhancements by integrating an illumination model and a noise model tailored for different scenes, which consider the unique interplay of haze, lighting, and sensor constraints.
Algorithm 1 Training pipeline of MToIE
@@
6:Trained network parameters $\theta_{\text{MToIE}}$ for MToIE model
@@
19:         Update network parameters $\theta_{\text{MToIE}}$
@@
III MToIE: Multi-Task-oriented Imaging Enhancer
III-A Network Architecture
Fig. 2 shows MToIE integrates a residual block-driven encoder-decoder framework with multi-task-oriented node learning (TNL) to handle three specific tasks: ID, LLIE and NHIE. To further enhance restoration performance, the framework incorporates SA and MRFE modules, strengthening the network’s focus on critical regions and features, ensuring robust handling of complex nighttime imaging scenarios. Finally, the network’s optimization is guided by a hybrid loss function, which balances multiple objectives, including image reconstruction quality and feature consistency, ensuring improved performance under challenging imaging conditions.
III-B Residual learning block
The residual learning block (RLB) is the foundational component of MToIE, primarily used for feature extraction and inference. As shown in Fig. 3, it comprises three key components: convolution, normalization, and activation functions. The convolution operation, utilizing local receptive fields and $3\times 3$ kernels, extracts detailed features from the input data. Layer normalization ($\mathcal{LN}$) is used to reduce internal covariate shift, enhancing training stability. The parametric rectified linear unit (PReLU) $\sigma$ is used, with adaptive parameters that improve the model’s expressive capability. Mathematically, a convolutional layer ($\mathcal{C}$) can be expressed as
@@ III-B Residual learning block
III-C Task-oriented Node Learning
@@
3:function MToIE($\mathbf{F_{e}}$)
@@
III-C1 Task-Specific Module
@@
III-C2 Self Attention-guided Node Learning Module
@@
III-D Multi-Receptive Field Enhancement Module
@@
III-E Self Attention Module
@@
III-F Hybrid Loss Function
@@ III-F3 Total Loss
IV Experiments and Discussions
In this section, we conduct extensive experiments to highlight the outstanding ID, LLIE, and NHIE performance of MToIE. We first provide an overview of the datasets and explain the implementation details. To illustrate the effectiveness of MToIE, we then conduct both quantitative and qualitative analyses, comparing it with various state-of-the-art methods using synthetic and real-world low-visibility images. Additionally, we perform ablation studies to evaluate the contribution of each proposed module. Finally, we analyze and compare MToIE in the context of the saliency detection task.
IV-A Datasets and Implementation Details
IV-A1 Training and Testing Datasets
The training and testing datasets include realistic single image dehazing (RESIDE)-OTS [24], which incorporates depth information, and the composite degradation dataset (CDD) [25]. Based on the imaging model in Sec. II, we extract atmospheric light values according to real conditions to synthesize more realistic degraded images [21]. By creating degraded images that closely resemble real-world conditions, we address the limitation of insufficient paired training datasets, thus enhancing the network’s restoration performance. During training, paired images are segmented into $256\times 256$ patches to achieve data augmentation and improve the network’s generalization capability. To evaluate the inference capabilities of MToIE, we use a variety of synthetic and real-world low-visibility images.
TABLE I: Comparison of ID results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@ IV-A1 Training and Testing Datasets
MToIE 25.461 $\pm$ 5.966 0.945 $\pm$ 0.067 3.159 $\pm$ 0.781
TABLE II: Comparison of LLIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
MToIE 28.529 $\pm$ 4.448 0.964 $\pm$ 0.035 3.243 $\pm$ 0.807
TABLE III: Comparison of NHIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
MToIE 25.617 $\pm$ 3.423 0.927 $\pm$ 0.049 3.425 $\pm$ 0.846
IV-A2 Evaluation Metrics
IV-A3 Competitive Methods
To evaluate the effectiveness of the MToIE, we conducted quantitative and qualitative comparative analysis with competing methods, including ID methods DCP [9], FFANet [15], SDD [8], CEEF [26], ROP+ [10], DehazeFormer [16], and AoSRNet [21]; LLIE methods LIME [7], SDD [8] , CEEF [26], EFINet [27], ROP+ [10], SMNet [14], and AoSRNet [21]; and NHIE methods IENHC [11], GMLC [12], MRP [13], OFSD [28], 4KDehazing [19], CEEF [26] and ECNHI [22]. To ensure fairness and objectivity in the experiment, all methods are based on the source code published by the authors, and multiple experimental comparisons are carried out under the identical conditions to minimize the impact of extraneous factors on the results.
IV-A4 Experiment Platform
The MToIE is trained for 200 epochs using 3183 images. The adaptive moment estimation optimizer is responsible for updating the network parameters. The initial learning rate for MToIE is set as $\epsilon=1\times 10^{-3}$ and is reduced by a factor of 10 at the 60th, 120th, and 180th epochs. The MToIE was trained and evaluated within the Python 3.7 environment using the PyTorch package with 2 Xeon Gold 37.5M Cache, 2.50 GHz @2.30GHz Processors and 4 Nvidia GeForce RTX 4090 GPUs.
IV-B Synthetic Degradation Analysis
IV-B1 Dehazing

### v2

# Multi-Knowledge-oriented Nighttime Haze Imaging Enhancer for Vision-driven Intelligent Systems

@@ Abstract
Salient object detection (SOD) plays a critical role in vision-driven intelligent systems (VIS), facilitating the detection and segmentation of key visual elements in an image. However, adverse imaging conditions such as haze during the day, low light, and haze at night severely degrade image quality, and complicating the SOD process. To address these challenges, we propose a multi-knowledge-oriented nighttime haze imaging enhancer (MKoIE), which integrates three tasks: daytime dehazing, low-light enhancement, and nighttime dehazing. The MKoIE incorporates two key innovative components: First, the network employs a task-oriented node learning mechanism to handle three specific degradation types: day-time haze, low light, and night-time haze conditions, with an embedded self-attention module enhancing its performance in nighttime imaging. In addition, multi-receptive field enhancement module that efficiently extracts multi-scale features through three parallel depthwise separable convolution branches with different dilation rates, capturing comprehensive spatial information with minimal computational overhead. To ensure optimal image reconstruction quality and visual characteristics, we suggest a hybrid loss function. Extensive experiments on different types of weather/imaging conditions illustrate that MKoIE surpasses existing methods, significantly enhancing the accuracy and reliability of vision systems across diverse imaging scenarios. The code is available at https://github.com/Ai-Chen-Lab/MKoIE.
keywords:
@@
Vision-driven Intelligent Systems , Adverse weather , Multi-knowledge learning , Nighttime haze , Image restoration
journal: Knowledge-Based Systems
[…]
[1]organization=School of Computer Science and Engineering, addressline=University of Electronic Science and Technology of China, city=Chengdu, postcode=611731, state=Sichuan, country=China
[…]
[2]organization=Department of Logistics and Maritime Studies, addressline=The Hong Kong Polytechnic University, postcode=999077, state=Hong Kong, country=China
[…]
[3]addressline=Chengdu Union Big Data Tech Inc., city=Chengdu, state=Sichuan, country=China
{highlights}
We propose a new multi-knowledge-oriented model for nighttime haze imaging.
We develop a network with task-specific nodes for feature extraction.
MKoIE enhances imaging in low-light, daytime haze, and nighttime haze conditions.
Extensive experiments show that the superiority of our MKoIE.
1 Introduction
@@
Salient object detection (SOD) focuses on detecting and segmenting the most visually prominent objects within vision-driven intelligent systems (VIS) [1, 2]. It is critical for enhancing automation and accuracy in VIS, particularly in applications like quality control and safety monitoring. However, as shown in Fig. 1, adverse imaging environmental conditions, such as low light, daytime or nighttime haze, can significantly affect the accuracy of SOD. Specifically, optical scattering from haze particles not only reduces contrast and sharpness but also distorts the spectral content of the captured scene [3]. Reduced illumination further exacerbates this issue by limiting signal strength, necessitating longer exposures or higher sensor gains, which result in additional electronic noise [4]. These environmental factors necessitate robust VIS that can effectively model and compensate for light attenuation, atmospheric scattering, and sensor-induced artifacts while maintaining measurement fidelity.
@@
Figure 1: The upper and lower sections illustrate daytime imaging scenarios, where varying haze densities impact object visibility. SOD operates on haze devices as a general-purpose preprocessing module, filtering salient data locally and transmitting it to the cloud for VIS.
@@
To address the challenges posed by harsh imaging conditions, we propose a multi-knowledge-oriented nighttime haze image enhancement framework (MKoIE), which be used to handle three types of degradation tasks: ID, LLIE, and NHIE. The framework integrates multi-knowledge learning mechanisms, self-attention (SA) module, and multi-receptive-field enhancement (MRFE) module to effectively enhance image quality under complex degradation scenarios. By employing task-specific node learning mechanisms and effectively using daytime haze, low-light, and nighttime haze datasets for training, MKoIE learns common visual patterns (such as edge contours and texture information) across different imaging conditions and specific degradation factors (such as brightness loss and noise enhancement in low-light environments, as well as contrast reduction and scattering blur induced by haze), efficiently models three distinct types of degradation. In addition, a hybrid loss function is designed to balance image reconstruction quality and visual features. The main contributions of this work are summarized as follows
  • 1.
@@
    We propose a multi-knowledge-oriented nighttime haze imaging enhancer (MKoIE) that incorporates task-specific node learning mechanisms and adopts shared and separated networks for feature-shared and specific feature extraction to effectively address low-light, daytime haze, and nighttime haze degradation scenarios.
  • 2.
@@
    The MKoIE integrates SA mechanism and MRFE module to extract multi-scale features. Additionally, a hybrid loss function is suggested to jointly optimize MKoIE for robust performance under challenging conditions.
  • 3.
@@
    Extensive experiments demonstrate that MKoIE significantly outperforms existing methods in image quality improvement and is a fast and lightweight model that achieves higher accuracy and reliability in subsequent high-level vision task.
@@
The remainder of this paper is organized as follows. Degradation models for nighttime haze are given in Section 3. MKoIE is detailedly described in Section 4. Experimental results and discussions are provided in Section 5. Conclusions are drawn in Section 6.
2 Related Work
Deep learning-based methods have achieved remarkable performance in ID, LLIE, and NHIE. For daytime image hazing, researchers have proposed a variety of deep neural network methods, including single-image dehazing networks [15, 16, 24, 25], multi-scale feature fusion networks [26, 27], and frameworks that combine deep learning with physical scattering models [21, 23]. Similarly, deep learning approaches have also led to notable advancements in LLIE [10, 14, 20, 28, 29, 30, 31, 32]. Nighttime dehazing remains more challenging, as low-light conditions often introduce noise and color distortion, while haze causes light scattering and contrast reduction, further complicating visibility restoration. To address these issues, researchers have explored methods such as multi-dimensional information fusion [33, 34], combined low-light image processing algorithms [11, 12, 22, 35], and generative adversarial network-based models [19, 36]. These recent works have not only enriched the research methods, but also yielded more effective solutions for practical applications [37].
Daytime Image Dehazing He et al. [9] introduced the influential Dark Channel Prior approach, and other early dehazing methods similarly rely on the physical atmospheric scattering model to estimate parameters such as atmospheric light, transmission maps, and scene radiance [38, 39]. However, these physics-based methods often struggle under complex lighting conditions or when scene depth varies significantly. Qin et al. [15] proposed FFANet that fuses multi-scale features to enhance detail recovery in hazy images, which exhibits limited adaptability to scenes with non-uniform haze distributions. Gong et al. [27] developed a two-stage network, which combines multi-scale feature fusion with an adaptive learning mechanism.But introduces additional computational overhead and demands a large amount of labeled training data. Yin et al. [23] integrated the atmospheric scattering model into a multi-stage dehazing framework, effectively blending physics-based and learning-based approaches. However, their method shows limited generalization in real-world scenarios and requires additional parameter tuning to handle different haze conditions.
Low-light Image Enhancement Traditional LLIE methods [6, 7, 40], such as Retinex-based algorithms, histogram equalization and gamma correction, aim to enhance contrast by redistributing pixel intensity, clarifying dark regions but often amplifying noise. Recent state-of-the-art LLIE strategies incorporate deep neural networks. For instance, URetinexNet [30] unrolls a Retinex-based optimization problem into a deep network guided by implicit prior regularization, featuring modules for data-dependent initialization and iterative optimization, enabling adaptive illumination adjustment. Retinexformer [32] integrates Retinex theory with a Transformer architecture in a one-stage model that simultaneously estimates illumination and corrects artifacts caused by overexposure. Hou et al. [20] proposed a diffusion-based framework with global structure-aware curvature regularization to enhance contrast and preserve details for structure-aware contrast enhancement. However, challenges such as noise amplification, color distortion, and susceptibility to overexposure persist, limiting real-world applicability.
Nighttime Image Dehazing Optical model-based methods [11, 12, 13, 41] extend daytime dehazing models by estimating atmospheric light and transmission maps, often incorporating preprocessing to handle nighttime-specific distortions. Li et al. [12] proposed a glow removal technique that separates halo artifacts from the scene using layer decomposition and adjusts for multi-colored light sources. Koo et al. [36] introduced a GAN-based framework that decomposes images into haze and glow components, allowing the generator to specifically suppress halos from bright light sources. Zheng et al. [19] employed a channel-spatial mixer for depth estimation and a domain adaptation module to translate synthetic hazy data to real-world haze. Cong et al. [34] proposed a spatial-frequency interaction module to address localized distortions while preserving realistic brightness levels. Despite these advances, nighttime dehazing remains challenging due to the combined effects of dense haze, low illumination, and glare from artificial light sources. Even state-of-the-art methods often struggle with residual glow artifacts or color distortions in such conditions, indicating the need for more robust and adaptive solutions.
3 Physical Imaging Models
@@
It needs to be noted that in practical applications of VIS, Eq. 1 needs to be adjusted according to specific circumstances. Due to the complexity of light source, particularly at night, directly applying above formula may not be sufficiently accurate. The challenges posed by varying lighting conditions, haze, and sensor noise require a more customized approach. To address these issues, our method focuses on three key image restoration tasks in VIS: ID, LLIE, and NHIE. To improve model accuracy under these conditions, we propose enhancements by integrating an illumination model and a noise model tailored for different scenes, which consider the unique interplay of haze, lighting, and sensor constraints.
4 MKoIE: Multi-Knowledge-oriented Imaging Enhancer
4.1 Network Architecture
Figure 2: Overall architecture of the multi-knowledge-oriented nighttime haze image enhancement (MKoIE) framework. The framework is designed to simultaneously handle three types of degradation tasks: ID, LLIE, and NHIE. By integrating multi-knowledge learning mechanisms, SA module, and MRFE module, it effectively improves image quality under complex degradation scenarios. The whole process of MKoIE is summarized in Algorithm 1.
Algorithm 1 Training pipeline of MKoIE
@@ 4.1 Network Architecture
6:Trained network parameters $\theta_{\text{MKoIE}}$ for MKoIE model
@@
19:         Update network parameters $\theta_{\text{MKoIE}}$
@@
Fig. 2 shows MKoIE integrates a residual block-driven encoder-decoder framework with multi-knowledge-oriented node learning (TNL) to handle three specific tasks: ID, LLIE and NHIE. To further enhance restoration performance, the framework incorporates SA and MRFE modules, strengthening the network’s focus on critical regions and features, ensuring robust handling of complex nighttime imaging scenarios. Finally, the network’s optimization is guided by a hybrid loss function, which balances multiple objectives, including image reconstruction quality and feature consistency, ensuring improved performance under challenging imaging conditions.
4.2 Residual learning block
The residual learning block (RLB) is the foundational component of MKoIE, primarily used for feature extraction and inference. As shown in Fig. 3, it comprises three key components: convolution, normalization, and activation functions. The convolution operation, utilizing local receptive fields and $3\times 3$ kernels, extracts detailed features from the input data. Layer normalization ($\mathcal{LN}$) is used to reduce internal covariate shift, enhancing training stability. The parametric rectified linear unit (PReLU) $\sigma$ is used, with adaptive parameters that improve the model’s expressive capability. Mathematically, a convolutional layer ($\mathcal{C}$) can be expressed as
@@ 4.2 Residual learning block
4.3 Task-oriented Node Learning
@@
3:function MKoIE($\mathbf{F_{e}}$)
@@
4.3.1 Task-Specific Module
@@
4.3.2 Self Attention-guided Node Learning Module
@@
4.4 Multi-Receptive Field Enhancement Module
@@
4.5 Self Attention Module
@@
4.6 Hybrid Loss Function
@@ 4.6.3 Total Loss
5 Experiments and Discussions
In this section, we conduct extensive experiments to highlight the outstanding ID, LLIE, and NHIE performance of MKoIE. We first provide an overview of the datasets and explain the implementation details. To illustrate the effectiveness of MKoIE, we then conduct both quantitative and qualitative analyses, comparing it with various state-of-the-art methods using synthetic and real-world low-visibility images. Additionally, we perform ablation studies to evaluate the contribution of each proposed module. Finally, we analyze and compare MKoIE in the context of the saliency detection task.
5.1 Datasets and Implementation Details
5.1.1 Training and Testing Datasets
The training and testing datasets include realistic single image dehazing (RESIDE)-OTS [42], which incorporates depth information, and the composite degradation dataset (CDD) [43]. Based on the imaging model in Sec. 3, we extract atmospheric light values according to real conditions to synthesize more realistic degraded images [21]. By creating degraded images that closely resemble real-world conditions, we address the limitation of insufficient paired training datasets, thus enhancing the network’s restoration performance. During training, paired images are segmented into $256\times 256$ patches to achieve data augmentation and improve the network’s generalization capability. To evaluate the inference capabilities of MKoIE, we use a variety of synthetic and real-world low-visibility images.
5.1.2 Evaluation Metrics
5.1.3 Competitive Methods
To evaluate the effectiveness of the MKoIE, we conducted quantitative and qualitative comparative analysis with competing methods, including ID methods DCP [9], FFANet [15], SDD [8], CEEF [29], ROP+ [10], DehazeFormer [16], and AoSRNet [21]; LLIE methods LIME [7], SDD [8] , CEEF [29], EFINet [31], ROP+ [10], SMNet [14], and AoSRNet [21]; and NHIE methods IENHC [11], GMLC [12], MRP [13], OFSD [46], 4KDehazing [19], CEEF [29] and ECNHI [22]. To ensure fairness and objectivity in the experiment, all methods are based on the source code published by the authors, and multiple experimental comparisons are carried out under the identical conditions to minimize the impact of extraneous factors on the results.
5.1.4 Experiment Platform
The MKoIE is trained for 200 epochs using 3183 images. The adaptive moment estimation optimizer is responsible for updating the network parameters. The initial learning rate for MKoIE is set as $\epsilon=1\times 10^{-3}$ and is reduced by a factor of 10 at the 60th, 120th, and 180th epochs. The MKoIE was trained and evaluated within the Python 3.7 environment using the PyTorch package with 2 Xeon Gold 37.5M Cache, 2.50 GHz @2.30GHz Processors and 4 Nvidia GeForce RTX 4090 GPUs.
Table 1: Comparison of ID results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@ 5.1.4 Experiment Platform
MKoIE 25.461 $\pm$ 5.966 0.945 $\pm$ 0.067 3.159 $\pm$ 0.781
Table 2: Comparison of LLIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
MKoIE 28.529 $\pm$ 4.448 0.964 $\pm$ 0.035 3.243 $\pm$ 0.807
Table 3: Comparison of NHIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
MKoIE 25.617 $\pm$ 3.423 0.927 $\pm$ 0.049 3.425 $\pm$ 0.846
Figure 6: Qualitative comparisons with the SOTA NHIE methods on the CDD datasets. Below each image is its RGB histogram distribution.
Figure 7: isual comparisons of detail enhancement.
5.2 Synthetic Degradation Analysis
5.2.1 Dehazing